OpenStudy (anonymous):

Please Help Arrange the equations in the correct sequence to find the inverse of f(x)=y=x-4/33-x

OpenStudy (anonymous):

@Hoslos

OpenStudy (anonymous):

THw first step is to replace x with y. Second step is to make y the subject of the formula.

OpenStudy (anonymous):

Tiles: Sequence: 33-xy=y-4 -> 33x+4=y(1+x) -> 33x-xy=y+4 -> 33x +4 = y+xy -> y=f^(-1)(x)= 1+x/ 33x+4 -> 33x- xy=y+4 -> x=y-4/33-y ->

OpenStudy (anonymous):

You made a mistake in the beginning, where it was supposed to be 33y not just 33.

OpenStudy (anonymous):

Thats how it is on my test /:

OpenStudy (anonymous):

\[x=\frac{ y-4 }{ 33- y}\rightarrow y-4=33x-xy\] \[y+yx=33x+4\rightarrow y(1+x)=33x+4\]\[y=f ^{-1}(x)=\frac{ 33x+4 }{ 1+x }\]

OpenStudy (anonymous):

So, as I said, I initially replaced y with x and made y the subject.

OpenStudy (anonymous):

Okay thank you, but im missing the last one that goes in the sequence /:

OpenStudy (anonymous):

You mean the whole sequence you post it or the last line of your sequence?

OpenStudy (anonymous):

The last line

OpenStudy (anonymous):

Its okay you tried helping me I appreciate it!(: @Hoslos

OpenStudy (anonymous):

I know that , but I also want to make sure you do not hold doubts in your brain. I know the answer for the last line though. It is the substitution of x into y. From which if you make y the subject, you shall get the answer of inverse, which was the one I found.

OpenStudy (anonymous):

Hope that satisfied.

OpenStudy (anonymous):

Thank you! and yes it does help a lot!!!(: @Hoslos

OpenStudy (anonymous):

Glad to hear.